OpenStudy (anonymous):

Co-ordinates= (3 ,2.5) and (-1.5 ,4). Work out m from y=mx+c and rearrange into that formula. I know I have to use these two equations: m=(y2-y1) / (x2-x1) AND y-y1= m(x-x1) Co-ordinates= (3 ,2.5) and (-1.5 ,4). Work out m from y=mx+c and rearrange into that formula. I know I have to use these two equations: m=(y2-y1) / (x2-x1) AND y-y1= m(x-x1) @Mathematics

OpenStudy (perl):

you can use y = mx + b for both

OpenStudy (perl):

m = ( 4- 2.5) / ( -1.5 - 3)

OpenStudy (mimi_x3):

If you know then substitute it in. x1 = 3 , y1 = 2.5 , x2 = -1.5 , y2 = 4

OpenStudy (perl):

m = ( 4- 2.5) / ( -1.5 - 3) = -1/3 so we have y = -1/3 x + b , now we have to find b

OpenStudy (perl):

plug in either point, i will use (3,2.5) 2.5 = -1/3 ( 3) + b , solve for b

OpenStudy (perl):

so b = 7/2 y = -1/3 x + 7/2

OpenStudy (anonymous):

Ah, I just realised where I went wrong. I wrote down 2 instead of 1.5 for 4-2.5. Thanks perl and Mimi.

OpenStudy (mimi_x3):

lols, i didnt really help xD

OpenStudy (perl):

yah mim, you didnt really help

OpenStudy (anonymous):

you 'assisted' then XD

OpenStudy (mimi_x3):

lols, its MIMI !

OpenStudy (anonymous):

Hang on, isn't it -0.33 recurring, rather than -1.3?